OpenStudy (anonymous):

Jerry receives a 15% employee discount at the store where he works. If he pays $168.30 for an item, what was the original price of the item

OpenStudy (mathmate):

Hints: Discounted price = list price *(1-discount) Example: John paid 175 for a printer at 30% discount. What is the list price. 175=List*(1-0.30) Solve for List List =175/0.7=250 Answer: The original (list) price was $250.